From GlueXWiki
Jump to: navigation, search

April 14, 2016 Drift Chamber meeting


  1. Instructions for Bluejeans meeting connection
  2. Meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. Detector updates
  2. Electronics (Fernando, Chris, Nick)
  3. NIM Papers
  4. fADC125 meeting (Naomi, Cody, Beni)
  5. Other


Participants: Curtis, Naomi, Mike (CMU) Eugene, Fernando, Beni, Luke, Simon, Nick, Chris, Vlad, Sergey, Amanda (detector group), Tyler (detector group), Pablo (detector group), and Lubomir (JLab).

Detector updates

- CDC has been successfully fixed after the "near-miss" event. The card that Nick suspected for noise was tested with a scope and was OK. Fernando: "the electronics saved the day" (i.e. the detector).

- Vlad reported on the FDC gain calibration. He did this before, but now the cathode resolution was improved and he uses now higher statistics (~1M hits per chamber). In the attached plots we see the distribution of the gain coefficients is getting narrower with each iteration - 4 in total. Some channels tend not to be calibrated producing tails in the distributions. Vlad will also work on parallelizing the calculations: separate job for each chamber.

- FDC monitoring: Lubomir proposed monitoring pages with the hit histograms, i.e. position and timing coincidence b/n the two cathode planes and the wire plane (example shown above). One page with 6 histograms per package. Mike: timing is difficult at early reconstruction, but position correlation is easy. Later Mike implemented such histograms. The complain was that this is next level of reconstruction and for the monitoring it is better to use just the occupancies.

- We discussed a possible photon background in that Hall based on measurements (logentry linked) with Xe, for which the x-section is 10-100 times higher than for Ar for photon energies in 10-100kev range (picture attached). The observed background difference with Ar/Xe may indicate presence of such photon background. More results and discussions will follow.

fADC125 (submitted by Naomi)

- Naomi is still looking at the differences between firmware and emulated output using the long mode (8) cosmic runs.

  • Run 11016 2 CDC fadcs were playing up: roc25 slot 9, even ch 36-40 and roc28 slot 16, even ch 36-40 - there were small errors in 4/5 of hits (firmware time too large by 1 or 2 units (0.8ns), and integral too small by 5 to 8). There were a small number of problems with the FDC fadcs (differences in time and pedestal) but nothing significant.
  • Run 11017 All CDC & FDC fadcs were ok. For this run the trigger holdoff had been increased to 30us.
  • Run 11075 There were many errors in the same two CDC fadc as in Run 11016, and also in one FDC fadc: roc62 slot 17 even ch 36-42 show same behaviour. For this run the bufferlevel had been set to 1.

- Naomi noticed that the fadc configuration parameters in the BOR data read back from the fadcs in run 11075 is corrupt for these 3 boards, ch 36-41 (all) in the CDC boards, and channels 36-47 in FDC roc 62. These include more than the dodgy channels mentioned above. This indicates a problem reading from the chip. If it were a write problem, the discrepancies between firmware/emulator would be much larger than observed. In run 11016 only the CDC boards had BORconfig errors, and run 11017's BOR decoding crashed.